OpenStudy (anonymous):

Your piggy bank has a total of 20 coins in it; some are dimes and some are nickels. If you have a total of $1.25, how many nickels do you have? 15 10 5 20

OpenStudy (anonymous):

d + n = 20 --> d = 20 - n .10d + .05n = 1.25 Now just sub 20 - n in for d in the 2nd equation :)

OpenStudy (anonymous):

still dont get it

OpenStudy (anonymous):

d stands for dimes n stands for nickels your question says d + n = 20 (there are 20 coins consisting of dimes and nickels) .10d + .05n = 1.25 (dimes are worth .10 and nickels are worth .05 and there is a total of 1.25) do you understand this so far ?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

now we have 2 equations d + n = 20 .10d + .05n = 1.25 change the first equation to d = 20 - n because you are looking for n. now sub 20 - n in for d in the other equation. .10(20 - n) + .05n = 1.25 now distribute the .10 through the parenthesis. Can you do that ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

show me :)

OpenStudy (anonymous):

thats what she said lol okay hold on

OpenStudy (anonymous):

lol....get your mind out of the gutter